OpenStudy (anonymous):

given that 4k-2, 18, 9k-1 are consecutive terms is a arithmetic sequence find K??

OpenStudy (cwrw238):

18 - (4k-2) = d (common difference of the AS) also 9k - 1 - 18 = d so 18 - 4k + 2 = 9k - 19 solve for k